OpenStudy (anonymous):

What is the value of 54 – 4a2 + 3b3 when a = −2 and b = 4?

OpenStudy (anonymous):

@ispike plz help

OpenStudy (anonymous):

plz explain

OpenStudy (michele_laino):

please if I substitute a=-2 and b=4, into your expression, i get this: \[54-4(-2)^{2}+3(4)^{3}=54-4*(4)+3*(64)=...\]

OpenStudy (anonymous):

ok give me sec to figure it out plz

OpenStudy (anonymous):

i will give my my answer in a second

OpenStudy (anonymous):

ok so far i have 54-16+192

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

154?

OpenStudy (michele_laino):

54-16+192=38+192=...

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

i did the PEMDAS thing

OpenStudy (anonymous):

you add before you subtract

OpenStudy (michele_laino):

\[54-16+192=(54-16)+192=38+192=...\]

OpenStudy (anonymous):

i got 230

OpenStudy (michele_laino):

that's right! because subtraction and addition has the same precedence, so you have to solve for first the operand which you encounter for first in going from left to right!